OpenStudy (anonymous):

margo's monthly payment on a loan is $1000. To refinance the loan, it will $1500 in fees and $850 each month. after how many months will the new loan save her money?

OpenStudy (amistre64):

after how many months will this save money eh...

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

the first month cost her a total of 1500+850 = 2350 dollar; which is 1350 more than usual; and she is saving 150 each month;

OpenStudy (amistre64):

1350/150 = 135/15 = 27/3 = 9 9 months